OpenStudy (anonymous):

25x/89w = 26b^R x and y equals ?

OpenStudy (lgbasallote):

seems impossible since you dont have a y...might want to recheck the question....

OpenStudy (anonymous):

thats exactly how it was written ?

OpenStudy (anonymous):

So the question is \[\frac{25x}{89w} = 26b^R\]? Find x and y?

OpenStudy (lgbasallote):

how about other information preceding/succeeding this one?

OpenStudy (anonymous):

sorry i meant w not y ?

OpenStudy (anonymous):

\[\frac{25x}{89w} = 26b^R\]\[25x=2314b^R\]\[x=92.56b^R\]

OpenStudy (anonymous):

\[\frac{25x}{89w} = 26b^R\]\[89w=\frac{25x}{26b^R}\]\[w=\frac{25x}{2314b^R}\]